OpenStudy (anonymous):

Carla walked 5/6 of a mile last weekend. This weekend she walked 1/3 as far. How far did Carla walk this weekend?

OpenStudy (anonymous):

Know it?

OpenStudy (anonymous):

5/6 of 1/3 = 5/6 x 1/3= 5/18, thank u